OpenStudy (anonymous):

SImplify: x+3/9/x+12/6 Really stuck on this problem

OpenStudy (campbell_st):

when dividing by a fraction, flip and multiply \[\frac{\frac{(x+3)}{9}}{\frac{(x+12)}{6}} = \frac{x+3}{9} \times \frac{6}{x+12}\] I'll leave you to simplify the fraction.

OpenStudy (anonymous):

See thats what i am having trouble on

OpenStudy (campbell_st):

\[\frac{6(x +3)}{9(x +12)}\] just simplify 6/9 the rest won't change

OpenStudy (anonymous):

2(x+3)/3(9x+12)

OpenStudy (campbell_st):

thats it... well done

OpenStudy (anonymous):

without the 9 oops lol thanks!

OpenStudy (campbell_st):

correct